### 01 May 2026 v3

# An Explicit Solution to Black–Scholes Implied Volatility

Abstract
This paper observes that the Black–Scholes call price can be written as the survival probability of an inverse Gaussian distribution, equivalently as a probability in variance space. Inverting this representation yields an analytically explicit formula for implied volatility in terms of the corresponding inverse Gaussian quantile function, with volatility on the left-hand side and only observable option inputs on the right-hand side. Numerical tests recover implied volatility to machine precision and, in a controlled setting, show the formula to be faster than a state-of-the-art benchmark.
Keywords: option markets, implied volatility, Black–Scholes, inverse Gaussian
JEL Classification: G13; C58
1 The Insight
The Black–Scholes formula is explicit from volatility to option price. Market practice, however, usually needs the reverse map, from an observed option price to the implied volatility that reproduces it. Since Black and Scholes (1973), implied volatility has become the standard convention for quoting and comparing option prices across strikes, maturities, assets, and dates. It is used, for example, throughout derivatives markets, empirical asset pricing, volatility forecasting, risk management, volatility-surface modelling, option-return studies, and tests of volatility risk premia. A formula for implied volatility therefore concerns one of the standard coordinates in which option markets are measured. Despite this importance, implied volatility is usually obtained by numerically inverting the Black–Scholes pricing formula. Given an option price, one searches for the volatility $\sigma$ for which the Black–Scholes price matches the observed price. The literature has therefore developed explicit approximations and asymptotic formulas such as Li (2008), analytical upper and lower bounds (Tehranchi, 2016), highly efficient iterative algorithms with Jäckel (2015, 2024) as a state-of-the-art benchmark, and exact but implicit representations through infinite series (Cui et al., 2021). Gerhold (2012) showed that implied volatility is not $D$-finite, ruling out a broad class of usual closed-form representations. The representation derived below is explicit in a different sense. It does not express implied volatility through elementary or $D$-finite functions, but reduces the Black–Scholes inversion exactly to the evaluation of a specified inverse Gaussian quantile.
For a call option with price $C$, strike $K$, forward $F$, risk-free discount factor $D$, and maturity $T$, the result for $K\neq F$ is
$\sigma(K,C)=\frac{2}{\sqrt{T}}\left[\mathcal{F}^{-1}_{IG}\left((1-c)/m;\,\tfrac{2}{|k|},\,1\right)\right]^{-1/2},\qquad m=\begin{cases}1&K>F,\\ K/F&K<F\end{cases}$ (1)
where $c:=C/(DF)$ is the normalized call price, $k:=\log(K/F)$ is forward log-moneyness, and $\mathcal{F}^{-1}_{IG}$ denotes the inverse Gaussian quantile function. At the forward strike, the equation collapses to
$\sigma(K=F,C)=\frac{2}{\sqrt{T}}\Phi^{-1}\left(\frac{c+1}{2}\right)$ (2)
The link to the spot price is $S=DFe^{\delta T}$ with continuous dividend yield $\delta$.
The formula turns the conversion from price to implied volatility into an inverse Gaussian quantile evaluation. It places implied volatility on the left-hand side and observable option inputs on the right-hand side. The only non-elementary operation is this quantile function. Thus, the formula removes the outer Black–Scholes root search and shifts the numerical task to the stable and accurate evaluation of that quantile. The underlying distribution is well studied, and the required quantile is available in standard numerical computing libraries. In the numerical tests below, the formula recovers implied volatility to machine precision. It is also about $3.4$ times faster than Jäckel (2024)’s implementation, a high-accuracy benchmark that is itself known for extremely fast Black–Scholes implied-volatility inversion (see, e.g., Matić et al., 2020; vollib, 2023).
The identity also gives a probabilistic interpretation. The normalized call price is the probability that a Brownian motion with drift $k/2$ has not yet reached level $1$ by the clock time $4/v^{2}$. Equivalently, total implied variance is the corresponding generalized-inverse Gaussian quantile. Thus, implied volatility is not merely the output of a volatility root search, but can also be read as a distributional transform of the market option price.
2 Derivation
Let $v=\sigma\sqrt{T}$ be the total implied volatility. The normalized Black–Scholes European call price is
$c_{BS}(k,v)=\Phi\left(-\frac{k}{v}+\frac{v}{2}\right)-e^{k}\Phi\left(-\frac{k}{v}-\frac{v}{2}\right).$ (3)
The case $k=0$ is handled by Eq. (2). It remains to consider $k\neq 0$.
First take an out-of-the-money call, so $k>0$. Write $\mathcal{F}_{IG}(\cdot;\mu,\lambda)$ for the inverse Gaussian distribution function with mean parameter $\mu$ and shape parameter $\lambda$, following the standard parametrization in Chhikara and Folks (1989). Its survival function is
$1-\mathcal{F}_{IG}(x;\mu,\lambda)=\Phi\left(-\frac{\sqrt{\lambda x}}{\mu}+\frac{\sqrt{\lambda}}{\sqrt{x}}\right)-e^{2\lambda/\mu}\Phi\left(-\frac{\sqrt{\lambda x}}{\mu}-\frac{\sqrt{\lambda}}{\sqrt{x}}\right).$ (4)
With $\lambda=1$, $\mu=2/k$, and $x=4/v^{2}$, this becomes
$c_{BS}(k,v)=1-\mathcal{F}_{IG}\left(\frac{4}{v^{2}};\frac{2}{k},1\right),\qquad k>0.$ (5)
Inverting Eq. (5) gives
$v(k,c)=\frac{2}{\sqrt{\mathcal{F}^{-1}_{IG}\left(1-c;\frac{2}{k},1\right)}},\qquad k>0.$ (6)
Finally, $\sigma=v/\sqrt{T}$. For $k<0$, the call is in the money. Its out-of-the-money counterpart has log-moneyness |k|=k|k|=-k and normalized call price
$\tilde{c}=1+\frac{F}{K}(c-1),\qquad 1-\tilde{c}=(1-c)\frac{F}{K}.$ (7)
Combining this transformation with Eq. (6) gives Eq. (1), the analytically explicit Black–Scholes implied-volatility representation.
The corresponding put formula follows from put–call parity,
$c-p=1-\frac{K}{F}=1-e^{k}.$ (8)
For a put with normalized price $p=P/(DF)$ and $K\neq F$, the probability argument in the inverse Gaussian quantile is $(e^{k}-p)/m$. Hence
$\sigma(K,P)=\frac{2}{\sqrt{T}}\left[\mathcal{F}^{-1}_{IG}\left((e^{k}-p)/m;\tfrac{2}{|k|},1\right)\right]^{-1/2},\qquad m=\begin{cases}1&K>F,\\ K/F&K<F.\end{cases}$ (9)
Probabilistic interpretation.
The identity also gives a direct probabilistic interpretation of the call price itself, analogous in spirit to the familiar probability reading of Black–Scholes delta. For an out-of-the-money call with $k>0$,
$c_{BS}(k,v)=\mathbb{P}\left(Y_{k}>\frac{4}{v^{2}}\right),\qquad Y_{k}\sim IG\left(\frac{2}{k},1\right).$ (10)
This representation has a first-passage-time interpretation. Let $B_{t}$ denote standard Brownian motion and define
$\tau_{k}:=\inf\left\{t>0:B_{t}+\frac{k}{2}t=1\right\}.$ (11)
Then $\tau_{k}$ is the time it takes a Brownian motion with drift $k/2$ and diffusion volatility $1$ to reach level $1$. For this drift and boundary, $\tau_{k}\sim IG\left(\frac{2}{k},1\right)$, so $Y_{k}$ is identical in distribution to $\tau_{k}$. Therefore, the normalized call price can also be interpreted as the probability that this process has not yet reached level $1$ by the variance-implied clock time $4/v^{2}$.
Equivalently, with $Z_{k}:=4/Y_{k}$,
$c_{BS}(k,v)=\mathbb{P}\left(Z_{k}<v^{2}\right),\qquad Z_{k}\sim GIG\left(\frac{1}{2},\frac{1}{4},k^{2}\right),$ (12)
where $GIG$ denotes the generalized inverse Gaussian distribution (see, e.g., Jørgensen, 1982). This form moves the probability statement into variance space. Since $k$ can be thought of as the log-return required to reach the strike, $Z_{k}$ is the corresponding variance-space hurdle. The normalized call price is therefore the probability that implied variance $v^{2}$ clears this moneyness-induced barrier. Thus, total implied variance is identified as a quantile of a generalized inverse Gaussian distribution,
$v(k,c)^{2}=\mathcal{F}_{GIG}^{-1}\left(c;\tfrac{1}{2},\tfrac{1}{4},k^{2}\right),\qquad k>0.$ (13)
This differs from return-space probability readings. For example, delta can be viewed as the probability that the terminal return clears the strike hurdle. Carr and Madan (2009) show that call prices can also be given a return-space probability interpretation under a suitable change of measure. The representation here is different because it assigns the probability statement to variance space rather than return space. The additional insight is that the market option price is not only a value obtained by averaging terminal payoffs, but also a probability level on a moneyness-specific variance scale. Implied volatility is then obtained by reading off the corresponding quantile on that scale.
3 Numerical Evaluation
I evaluate the formula on a deterministic grid of total implied volatilities and Black–Scholes call deltas. Total volatility is set to $v\in\{0.01,0.05,0.10,\ldots,2.00\}$, and call delta is set to $\Delta\in\{0.05,0.20,0.30,0.45,0.55,0.70,0.80,0.95\}$, giving 328 test cases. For each pair $(v,\Delta)$, the corresponding forward log-moneyness is obtained from the Black–Scholes delta relation,
$k=v\left(\frac{v}{2}-\Phi^{-1}(\Delta)\right).$ (14)
The normalized call price is then generated from Eq. (3). The implied volatility formula is applied to this price, and the recovered total volatility is compared with the original input value.
Figure 1: Recovery of total implied volatility $v$ from the inverse Gaussian quantile formula (Explicit) and scalar speed comparison against Jäckel (2024)’s benchmark.
As benchmark I use the reference implementation of Jäckel’s Let’s Be Rational, downloaded from the author’s website (Jäckel, 2024). This is a stringent comparison, since the algorithm is designed to recover implied volatility to maximum attainable double-precision accuracy in as little as two iterations. Both methods recover the original total volatility to machine precision. The inverse Gaussian quantile formula attains a mean absolute recovery error of $2.24\times 10^{-16}$, compared with $2.12\times 10^{-16}$ for Jäckel’s implementation.
I also report a simple scalar timing comparison. This timing exercise is meant as an indicative implementation check, not as a general performance claim. The timing results should therefore be read as implementation-specific rather than as a general ranking of algorithms. One timing run consists of evaluating the 328 test cases 5,000 times. This corresponds to 1,640,000 implied-volatility evaluations per method. Timings are recorded over five such runs. The timing experiment was run on an HP EliteBook 840 G8 with an 11th Gen Intel Core i5-1145G7 processor at 2.60 GHz, using 4 physical cores and 8 logical processors, under Microsoft Windows. The experiment was implemented in Python, but both computational routines were evaluated in native compiled code. Table 1 reports the speed and recovery-error summary, while Figure 1 shows the recovery errors and timing variation across runs. On this benchmark grid and implementation, the inverse Gaussian quantile formula requires $0.305$ microseconds per evaluation, compared with $1.038$ microseconds for Jäckel’s implementation.
Table 1: Speed and accuracy comparison on the delta grid. Speed is measured on 1,640,000 implied-volatility evaluations per timing run.
Speed Absolute recovery error
total seconds microseconds/eval. mean max
Explicit 0.500 0.305 $2.24\times 10^{-16}$ $1.33\times 10^{-15}$
Jäckel 1.702 1.038 $2\times 10^{-16} 1.22×1$$2\times 10^{-15}$
4 Conclusion
This paper gives a compact analytically explicit representation of Black–Scholes implied volatility from observed option inputs. The key step is to recognize the normalized call price as a survival probability of an inverse Gaussian distribution. Inverting this identity writes implied volatility as a reciprocal square root of the corresponding inverse Gaussian quantile, with implied volatility on the left-hand side and observable inputs on the right-hand side. The only non-elementary object is the quantile function, so the representation removes the outer Black–Scholes root search while shifting the remaining numerical task to stable and accurate quantile evaluation. Equivalently, in variance space, the normalized option price determines a probability level and total implied variance is the corresponding generalized inverse Gaussian quantile. The representation also gives a probabilistic interpretation. For fixed strike and forward, the normalized call price is the probability that a Brownian motion with drift $k/2$ has not yet reached level $1$ by the clock time $4/v^{2}$. Implied volatility can therefore be viewed not only as the volatility that makes Black–Scholes match the observed option price, but also as the value obtained from a moneyness-specific distributional quantile of the normalized option price. The numerical evidence confirms that the inverse Gaussian representation is directly computable. On the deterministic test grid, it recovers total implied volatility to machine precision, with a mean absolute recovery error of $2.24\times 10^{-16}$. In the particular scalar benchmark considered here, it requires $0.305$ microseconds per evaluation, compared with $1.038$ microseconds for Jäckel’s Let’s Be Rational implementation (Jäckel, 2015, 2024).
The result does not change the economic content of Black–Scholes. It clarifies its inverse. Implied volatility remains the volatility that makes Black–Scholes match the observed option price, but it can be represented analytically through a standard distributional quantile.